How do you solve Rational Exponent Equations such as:
27=x^3/2

Answer by stanbon(75887)

27=x^3/2
----
Raise both sides to the (2/3) power to get:
x = 27^(2/3)
-----
Comment: The "3" in the denominator tells you to take the cube root.
x = 3^2
----
Comment: The "2" in the numerator tell you to square the base.
x = 9
===========
Cheers,
Stan H.